Suppose that you have found the optimal risky combination using all risky assets available in the economy, and that this optimal risky portfolio has an expected return of 0.1 and standard deviation of 0.2. The T-bill rate is 0.05.

If your risk-return preferences are best described by the utility function in this class, with a risk-aversion coefficient of 5.2. What is the expected return on your optimal complete portfolio?

Round your answer to 4 decimal places. For example if your answer is 3.205%, then please write down 0.0321.
Answer #1

Proportion in optimal risky portfolio=(return on optimal risky portfolio-risk free rate)/(A*standard deviation of optimal risky portfolio^2)=(0.1-0.05)/(5.2*0.2*0.2)=0.24038

Expected return on optimal complete portfolio=0.24038*0.1+(1-0.24038)*0.05=0.06202